Question one

a)

i) ii)

Given the following data derive the consumption function ( 5 marks)

Derive the consumption function (4 marks)
Assuming that income (Y) is 1800, comment on the level of induced consumption (2 arks)

House hold Saving

-110

-90

-70

-50

Income (Y)

200

300

400

500

b) DemonstrateyourunderstandingonRatcheteffectusingawelllabeleddiagram(4mks)
Answer #1

♡ At income leve 4-200, Savey =-110 4 c=310 мрс Ас IMPS= = 2 20.20 MPC = I-MPS = 1-0.2=0.8 When Y=200 C+0.84 = 310 Ĉ = 310-16